# Can LLMs Help Uncover Insights about LLMs? A Large-Scale, Evolving Literature Analysis of Frontier LLMs

### v2

@@ 1 Introduction
In this study, we conduct the most extensive quantitative literature analysis of frontier LLMs to date. Our dataset, LlmEvalDB, comprises 18,127 experimental records extracted from 1,737 papers, along with 359 referenced papers describing datasets used in these experiments. LlmEvalDB can also be dynamically updated with minimal human effort to incorporate results from new models or studies as they are published. This comprehensive dataset enables us to present new insights into LLM performance, a first in the field where previous studies on extracting leaderboards from machine learning papers have primarily focused on improving data extraction accuracy (Kardas et al., 2020; Yun et al., 2024).
To achieve this, we experiment with an automated approach to data extraction that efficiently processes research literature. This approach uses an LLM to scan the arXiv database, identify relevant papers with experimental results on target models, and extract experimental results along with pertinent attributes. This reduces the manual effort needed for surveying and extraction by more than 93% compared to expert manual data extraction. The approach employs schema-driven extraction (Bai et al., 2023) and context augmentation from paper contents to capture essential attributes for a comprehensive understanding of evaluation results. Additionally, the process creates dataset descriptions that summarize key characteristics of the dataset associated with the evaluation record, enhancing the dataset’s potential utility in literature analysis.
Our analysis shows that LlmEvalDB can efficiently replicate a previous manual analysis (Sprague et al., 2024), which found that CoT reasoning primarily benefits mathematical and symbolic reasoning tasks. We also identify three novel insights from the dataset on prompting configurations. First, in-context learning (ICL) enhances coding & multimodal tasks but offers limited benefits for mathematical tasks compared to a zero-shot setup. Second, our qualitative analysis reveals specific characteristics of datasets that tend to reduce the effectiveness of CoT prompting and ICL. Third, we find that CoT prompting with demonstrations—commonly implemented through ICL—typically yields better performance than CoT without demonstrations. However, the relative improvement of CoT over standard direct prompting remains stable regardless of the number of demonstrations used (i.e., it holds true in both zero-shot and few-shot settings). We release our dataset and source code to support the ongoing automated integration of new findings in the literature.  [^1] The code is available at https://github.com/JJumSSu/meta-analysis-frontier-LLMs, and the dataset can be accessed at https://huggingface.co/datasets/jungsoopark/LLMEvalDB.
@@ 2.4 Dataset Description Generation
In addition to extracting structured representations of experiments, we also generate summary descriptions of the relevant tasks and datasets (Fig. 3) that can aid in-depth literature analysis, such as classifying records by research subarea. We use a two-stage approach to create these descriptions. Initially, the LLM generates descriptions using its internal knowledge based on the dataset name and subset, which is cost-effective. When the LLM is uncertain, especially with lesser-known datasets or due to a knowledge cut-off, we prompt it to extract descriptions using the full content of the source papers as a reference. The source papers can be either the paper containing the table or the original dataset paper cited within it, which is retrieved using the BibTex references obtained from section 2.3.
@@ 4.2 Which Tasks Benefit from ICL?
Figure 7 presents the results. In contrast to CoT’s strong performance in mathematical reasoning, ICL shows only modest benefits for math tasks. However, ICL demonstrates more substantial improvements in coding and multimodal applications despite considerable performance variability across different cases. This variance may be due to the broad scope of multimodal tasks, including speech and image processing. ICL shows more uniform improvements over zero-shot performance across different categories compared to the study in Fig. 5 (std: 8.3). Note that we did not account for the number of in-context examples used; we only distinguished between the use of in-context examples and zero-shot examples. We also plot the performance improvement distribution of more versus fewer in-context examples in appendix C and find a similar overall median. This suggests that the presence of demonstrations is more important than their quantity.
@@ 4.3 Which Dataset Characteristics Negatively Affect CoT and ICL?
The two approaches show different patterns in other aspects. CoT significantly degraded performance in faithfulness and factual verification tasks, possibly due to introduced reasoning biases (Shaikh et al., 2023; Chua et al., 2024). In contrast, demonstrations led to notable declines in affective analysis and structured prediction tasks, where examples may bias the model toward imitating specific patterns rather than genuine reasoning. This also partially aligns with the finding from Zhang et al. (2022) that demonstrations increase the confidence of model predictions on captured superficial patterns, possibly incurring biases for structure prediction tasks such as NER.
CoT Reasoning Median Q1 Q3
$\Delta$ (Few-shot, Zero-shot) 3.0 0.4 9.2
$\Delta$ (More-shot, Less-shot) 3.1 0.4 9.0
Table 5: Statistical distribution of performance improvements when using varying levels of demonstrations in CoT prompting. $\Delta$(A, B) refers to the performance improvement of A over B, Q1 and Q3 represent the first and third quartiles, respectively.
$\Delta$ (CoT, Direct Prompting) Median Q1 Q3
Few-shot 0.9 -1.2 3.7
Zero-shot 1.3 -0.4 4.7
Table 6: Statistical distribution of performance improvements for CoT compared to direct prompting given different levels of demonstrations.
@@ 4.4 How do CoT and ICL Interact?
Building on previous studies section 4.1 and section 4.2, our analysis uncovers how CoT reasoning interacts with demonstrations in prompting, highlighting their complementarity. We extract instances from LlmEvalDB where CoT is used with varying levels of demonstrations, as well as instances that compare CoT and direct prompting under different demonstration levels (zero-shot vs. few-shot), all while keeping other attributes constant.
@@
From Table 5, demonstrations consistently enhance CoT performance: few-shot CoT outperforms zero-shot CoT by 3.0 points, and increasing the number of demonstrations yields a similar 3.1-point improvement. These comparable gains suggest that the mere presence of demonstrations, rather than their quantity, drives the improvement. However, from Table 6, we can see that when comparing CoT against standard prompting, the benefits remain relatively constant regardless of demonstration use. CoT’s improvement over standard prompting is similar in both few-shot and zero-shot settings. This indicates that while demonstrations enhance both approaches, they do not amplify CoT’s relative advantage over standard prompting.
@@ 5 Related Work
Our study focuses on extracting experimental results of frontier proprietary LLMs (Achiam et al., 2023; Anthropic, 2024; Team et al., 2023), with a specific emphasis on target attributes that incorporate information about prompting methods (Brown et al., 2020; Wei et al., 2022; Min et al., 2022). In the context of prompting, prior studies have systematically analyzed the mechanisms behind prompting methods, focusing either on the use of in-context examples (Min et al., 2022; Lampinen et al., 2022; Weber et al., 2023; Zhang et al., 2022) or techniques that elicit reasoning, such as CoT prompting (Wei et al., 2022; Shaikh et al., 2023; Wang et al., 2023; Turpin et al., 2024; Sprague et al., 2024; Liu et al., 2024). Conversely, we examine the model’s behavior by conducting a literature analysis, which compiles data from scientific sources to reveal insights.
@@ Appendix A Details and Trials During Extraction
Initially, we asked the LLM to generate descriptions using only the dataset name and subset information from its internal knowledge. This approach was cost-effective as it didn’t require processing additional context. However, when the LLM was uncertain—particularly with lesser-known datasets or those beyond its knowledge cutoff—it acknowledged its uncertainty, prompting the second stage of extraction from source papers.
When LLM knowledge was insufficient, we directly extracted dataset descriptions from source papers. These source papers could be either the paper containing the table being processed or the original dataset paper cited by it. We used citation tags from section 2.3 and rule-based heuristics to link to external arXiv papers associated with the original dataset. We then used the full content of these source papers to prompt the LLM to extract the dataset descriptions again, using the paper’s contents as a reference.
@@ C.1 Which Categories Benefit from CoT?
CoT Reasoning Original Results Filtered Results
@@
$\Delta$ (Few-shot, Zero-shot) 3.0 0.4 9.2 5.0 2.3 10.4
$\Delta$ (More-shot, Less-shot) 3.1 0.4 9.0 5.0 2.3 10.7
Table 8: Statistical distribution of performance improvements when using varying levels of demonstrations in CoT prompting. Original Results refers to the results from Table 5 from the submission (without filtering any papers). Filtered Results refer to the results from the filtered papers that are published in top conferences. $\Delta$(A, B) refers to the performance improvement of A over B, Q1 and Q3 represent the first and third quartiles, respectively.
$\Delta$ (CoT, Direct Prompting) Original Results Filtered Results
Median Q1 Q3 Median Q1 Q3
Few-shot 0.9 -1.2 3.7 1.3 0.0 3.0
Zero-shot 1.3 -0.4 4.7 1.3 0.0 3.6
Table 9: Statistical distribution of performance improvements for CoT compared to direct prompting given different levels of demonstrations. Original Results refers to the results from Table 6 from the submission (without filtering any papers). Filtered Results refer to the results from the filtered papers that are published in top conferences.

### v3

@@ 1 Introduction
In this study, we conduct the most extensive quantitative literature analysis of frontier LLMs to date. Our dataset, LlmEvalDB, comprises 18,127 experimental records extracted from 1,737 papers, along with 359 referenced papers describing datasets used in these experiments. LlmEvalDB can also be dynamically updated with minimal human effort to incorporate results from new models or studies as they are published. This comprehensive dataset enables us to present new insights into LLM performance, where previous studies on extracting leaderboards from machine learning papers have primarily focused on improving data extraction accuracy (Kardas et al., 2020; Yun et al., 2024).
To achieve this, we experiment with an automated approach to data extraction that efficiently processes research literature. This approach uses an LLM to scan the arXiv database, identify relevant papers with experimental results on target models, and extract experimental results along with pertinent attributes. This reduces the manual effort needed for surveying and extraction by more than 93% compared to expert manual data extraction. The approach employs schema-driven extraction (Bai et al., 2023) and context augmentation from paper contents to capture essential attributes for a comprehensive understanding of evaluation results. Additionally, the process creates dataset descriptions that summarize key characteristics of the dataset associated with the evaluation record, enhancing LlmEvalDB’s potential utility in literature analysis.
Our analysis shows that LlmEvalDB can efficiently replicate a previous manual analysis (Sprague et al., 2024), which found that CoT reasoning primarily improves performance on mathematical and symbolic reasoning benchmarks. We also identify three novel insights from the dataset on prompting configurations. First, in-context learning (ICL) enhances coding & multimodal tasks but offers limited benefits for mathematical tasks compared to a zero-shot setup. Second, our qualitative analysis reveals specific characteristics of datasets that tend to reduce the effectiveness of CoT prompting and ICL. Third, we find that CoT prompting with demonstrations—commonly implemented through ICL—typically yields better performance than CoT without demonstrations. However, the relative improvement of CoT over standard direct prompting remains stable regardless of the number of demonstrations used (i.e., it holds true in both zero-shot and few-shot settings). We release our dataset and source code to support the ongoing automated integration of new findings in the literature.  [^1] The code is available at https://github.com/JJumSSu/meta-analysis-frontier-LLMs, and the dataset can be accessed at https://huggingface.co/datasets/jungsoopark/LLMEvalDB.
@@ 2.4 Dataset Description Generation
In addition to extracting structured representations of experiments, we also generate summary descriptions of the relevant tasks and datasets (Fig. 3) that can aid in-depth literature analysis, such as classifying records by research subarea. We use a two-stage approach to create these descriptions. Initially, the LLM generates descriptions using its internal knowledge based on the dataset name and subset, which is cost-effective. When the LLM is uncertain and refuses to answer, we prompt it to extract descriptions using the full content of the source papers as reference. The source papers can be either the paper containing the table or the original dataset paper cited within it, which is retrieved using the BibTex references obtained from section 2.3.
@@ 4.2 Which Tasks Benefit from ICL?
Figure 7 presents the results. In contrast to CoT’s strong performance in mathematical reasoning, ICL shows only modest benefits for math tasks. However, ICL demonstrates more substantial improvements in coding and multimodal applications despite considerable performance variability across different cases. This variance may be due to the broad scope of multimodal tasks, including speech and image processing. ICL shows more uniform improvements over zero-shot performance across different categories compared to the study in Fig. 5 (std: 8.3). Note that we did not account for the number of in-context examples used; we only distinguished between the use of in-context examples and zero-shot learning. We also plot the performance improvement distribution of more versus fewer in-context examples in appendix C and find a similar overall median. This suggests that the presence of demonstrations is more important than their quantity.
@@ 4.3 Which Dataset Characteristics Negatively Affect CoT and ICL?
The two approaches show different patterns in other aspects. CoT significantly degraded performance in faithfulness and factual verification tasks, possibly due to introduced reasoning biases (Shaikh et al., 2023; Chua et al., 2024). In contrast, demonstrations led to notable declines in affective analysis and structured prediction tasks, where examples may bias the model toward imitating specific patterns rather than genuine reasoning. This also partially aligns with the finding from Zhang et al. (2022a) that demonstrations increase the confidence of model predictions on captured superficial patterns, possibly incurring biases for structure prediction tasks such as NER.
Performance comparison Median Q1 Q3
Few-shot CoT -- Zero-shot CoT 3.0 0.4 9.2
Table 5: Effect of in-context demonstrations within Chain-of-Thought prompting.
Demonstration level $\Delta$(Median) $\Delta$(Q1) $\Delta$(Q3)
Zero-shot +1.3 --0.4 4.7
Few-shot +0.9 --1.2 3.7
Table 6: The performance comparison ($\Delta$) between CoT and standard prompting under matched demonstration settings. A positive $\Delta$ indicates that CoT outperforms standard prompting.
@@ 4.4 How do CoT and ICL Interact?
Building on previous work sections 4.1 and 4.2, we ask two related questions about the joint behavior of CoT and ICL:
  1. 1.
    Do demonstrations boost performance inside the CoT framework? We compare few-shot and zero-shot settings while keeping CoT prompting constant (Table 5), replicating the manual study from Wei et al. (2022) through our literature analysis.
  2. 2.
    Do demonstrations enable CoT to beat standard prompting? We compare CoT and standard prompting under the same demonstration level (Table 6).
@@
  1. 1.
    Demonstrations within CoT. Table 5 shows that adding demonstrations yields a median gain of +3.0 pts which is significant. This replicates the uplift reported by (Wei et al., 2022; Zhang et al., 2022b).
  2. 2.
    Demonstrations do not “rescue” CoT. Table 6 shows that, whether we include demonstrations or not, CoT fails to open a meaningful lead over standard prompting. The median gains are modest— +0.9 pts in the few-shot setting and +1.3 pts in zero-shot. In other words, demonstrations lift both prompting styles in parallel but do not turn CoT into the clearly better option.
@@ 5 Related Work
Our study focuses on extracting experimental results of frontier proprietary LLMs (Achiam et al., 2023; Anthropic, 2024; Team et al., 2023), with a specific emphasis on target attributes that incorporate information about prompting methods (Brown et al., 2020; Wei et al., 2022). In the context of prompting, prior studies have analyzed the mechanisms behind prompting methods, focusing either on the use of in-context examples (Min et al., 2022; Lampinen et al., 2022; Weber et al., 2023; Zhang et al., 2022a) or techniques that elicit reasoning, such as CoT prompting (Wei et al., 2022; Shaikh et al., 2023; Wang et al., 2023; Turpin et al., 2024). Conversely, we examine the model’s behavior by conducting a literature analysis, which compiles data from scientific sources to reveal insights.
@@ Appendix A Details and Trials During Extraction
Initially, we asked the LLM to generate descriptions using only the dataset name and subset information from its internal knowledge. This approach was cost-effective as it didn’t require processing additional context. However, the LLM may be uncertain—particularly with lesser-known datasets or those beyond its knowledge cutoff. To address this, we instructed the model to refuse to answer when unsure about dataset information (Bai et al., 2022). In such cases, a second stage is triggered, where the model is prompted to extract the necessary information directly from the source papers. When LLM knowledge was insufficient, we directly extracted dataset descriptions from source papers. These source papers could be either the paper containing the table being processed or the original dataset paper cited by it. We used citation tags from section 2.3 and rule-based heuristics to link to external arXiv papers associated with the original dataset. We then used the full content of these source papers to prompt the LLM to extract the dataset descriptions again, using the paper’s contents as a reference.
@@ C.1 Which Categories Benefit from CoT?
Original Filtered
Performance comparison Median Q1 Q3 Median Q1 Q3
Few-shot CoT -- Zero-shot CoT 3.0 0.4 9.2 5.0 2.3 10.4
Table 8: Effect of in-context demonstrations within Chain-of-Thought prompting. Original uses all papers; Filtered keeps only top-venue papers.
Demonstration level Original $\Delta$ Filtered $\Delta$
@@
Zero-shot +1.3 --0.4 +4.7 +1.3 0.0 +3.6
Few-shot +0.9 --1.2 +3.7 +1.3 0.0 +3.0
Table 9: CoT versus standard prompting when the demonstration setting is fixed. Positive $\Delta$ means CoT outperforms Direct prompting.Original $\Delta$ uses all papers; Filtered $\Delta$ keeps only top-venue papers.